# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ 3.2 Refined $\ell_{2}$ error characterization
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np},$ (5)
Theorem 2 is more refined compared to (3). First, it provides both upper and lower bounds for $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$. Second, there is no hidden constant in front of the leading term $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$. In addition, inspecting the proof of Theorem 2, we see that
and the same holds for $[\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})]^{1/2}$. Consequently, the right hand sides of both (5) and (6) are lower order terms compared to $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$ when $n,m\rightarrow\infty$. Indeed this recovers the naive $\ell_{2}$ bound (3) under appropriate sample size assumptions.
3.3 Uncertainty quantification
Now we move on to uncertainty quantification about the $\mathsf{RP\text{-}MLE}$ estimator $\widehat{\bm{\theta}}$ in order to perform, say, hypothesis testing on the item parameters $\bm{\theta}^{\star}$. Our starting point is the following characterization of the limiting distribution of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$. The analysis is deferred to Section 4.3 and the complete proof is deferred to Section E.
@@ 3.3 Uncertainty quantification
Instate the assumptions of Theorem 1 and fix $m$. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
$\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
Though simple, we record the following immediate consequence on the asymptotic normality of $\widehat{\theta}_{i}-\theta_{i}^{\star}$ for any single coordinate $i$.
Corollary 1.
Instate the assumptions of Theorem 1. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$ (7)
We compare our result with that in [CLOX23]. Theorem 8 therein studies the Rasch model and provides an inference result for a joint estimator of $\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}$. However, it requires a dense sampling scheme with $p\gtrsim 1/\sqrt{\log(n)/n}$ and $p\gtrsim n\log^{2}(n)/m^{2}$. It also requires that both $m$ and $n$ tend to infinity. These two assumptions are more restrictive than ours.
Theorem 3 is quite general as it allows various applications to inference. As an illustrative example, we demonstrate how to construct confidence intervals for a particular item parameter $\theta_{i}^{\star}$.
Fix any $\alpha>0$, the level of significance. Define
$\mathcal{C}_{i}^{+}(\alpha)$ $\coloneqq\widehat{\theta}_{i}+\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
$\mathcal{C}_{i}^{-}(\alpha)$ $\coloneqq\widehat{\theta}_{i}-\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
where $\Phi$ is the cdf for the standard normal distribution. Corollary 1 implies the following proposition.
@@
Suppose the statement in Theorem 3 holds. Then for any $(i,j)$, we have
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\geq\mathcal{C}_{i}^{+}(\alpha)\right]$ $=\alpha,$ (8a)
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\leq\mathcal{C}_{i}^{-}(\alpha)\right]$ $=\alpha,$ (8b)
and
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\right]=\alpha.$ (9)
This proposition can be extended to confidence intervals for any linear combination of $\bm{\theta}^{\star}$, such as $\theta_{i}^{\star}-\theta_{j}^{\star}$. These confidence intervals naturally enable hypothesis testings such as
$\mathcal{H}_{0}:\theta_{j}^{\star}\geq\theta_{i}^{\star},\qquad\mathcal{H}_{1}:\theta_{j}^{\star}<\theta_{i}^{\star}$
and
$\mathcal{H}_{0}:\text{item $i$ is a top-$K$ item},\qquad\mathcal{H}_{1}:\text{item $i$ is not a top-$K$ item}.$
Various other inference tasks have been considered in the ranking literature [LFL23][FLWY22, FLWY23]. We expect our general distributional characterization in Theorem 3 to be readily applied for these tasks.
4 Analysis

### v2

@@ 3.3 Asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and pseudo MLE
Instate the assumptions of Theorem 1. Consider $\mathrm{MRP\text{-}MLE}$ with $n_{\mathrm{split}}$ random splits with fixed $m$ and $p$. Suppose that the limits in (8) exist. Then as $n\rightarrow\infty$,
$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right]\left(\bm{H}^{\infty}\right)^{\dagger}\right).$ (10)
Theorem 3 provides an asymptotic result for inference of $\mathrm{MRP\text{-}MLE}$ when $m$ and $p$ are fixed and $n\rightarrow\infty$. In particular, it reveals the decrease in asymptotic covariance of $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ from $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ to $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ as $n_{\mathrm{split}}$ goes from $1$ to $\infty$. In what follows, we first connect this result with the asymptotic normality of $\mathrm{WP\text{-}MLE}$, which takes all possible item-item pairs with overlap instead of doing random pairing. Then we quantify the asymptotic variance of $\mathrm{MRP\text{-}MLE}$ in a special instance to see how much using multiple random splitting helps.
Asymptotic normality of weighted pseudo MLE.
Pseudo MLE (Zwi, 95) is a method for item parameter estimation similar to our approach $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$. Instead of random pairing, pseudo MLE forms all possible item-item pairs with overlaps. For the non-asymptotic analysis, the lack of independence between comparisons induced by the overlaps is clumsy. However, here we show that a variant of pseudo MLE is asymptotically normal and relates to our method $\mathrm{MRP\text{-}MLE}$.
Now we formally introduce the weighted pseudo MLE, denoted as $\mathrm{WP\text{-}MLE}$. Let $m_{t}$ be the total number of responses of user $t$ and $\widetilde{m}_{t}$ be the largest even number smaller or equal to $m_{t}$. Consider the negative log-likelihood functions
$\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j:\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right]$
and
$\mathcal{L}_{\mathrm{WP}}(\bm{\theta})\coloneqq\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
Then $\mathrm{WP\text{-}MLE}$ is defined as
$\widehat{\bm{\theta}}_{\mathrm{WP}}=\arg\min_{\bm{\theta}\in\mathbb{R}^{m},\bm{\theta}^{\top}\bm{1}_{m}=0}\mathcal{L}_{\mathrm{WP}}(\bm{\theta}).$ (11)
Similar to $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$, we also use the notation $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ when appropriate. The intuition behind this reweighting is to account for the different numbers of responses for each user. It can be easily shown that $\mathbb{E}_{\mathrm{r}}\mathcal{L}_{k}(\bm{\theta})=\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$. Similar to the result for $\mathrm{MRP\text{-}MLE}$, we have the following asymptotic normality result for $\mathrm{WP\text{-}MLE}$. The proof is deferred to Section D.2.
@@
Instate the assumptions of Theorem 1. Consider $\mathrm{WP\text{-}MLE}$ with fixed $m$ and $p$. Suppose that the limits in (8) exist. In addition, assume that for every fixed $\bm{\theta}$ obeying $\|\bm{\theta}-\bm{\theta}^{\star}\|_{\infty}\leq 10$, the following limit exists:
$\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
Then as $n\rightarrow\infty$,
$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}\left(\bm{H}^{\infty}\right)^{\dagger}\right).$ (12)
This theorem establishes the asymptotic normality of $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$. We observe that the asymptotic covariance of $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ is equal to the asymptotic covariance $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ as $n_{\mathrm{split}}$ tends to infinity. This connects $\mathrm{WP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$, showing that they are asymptotically equivalent in distribution when $n\rightarrow\infty$ and $n_{\mathrm{split}}\rightarrow\infty$.
Quantifying the shift of asymptotic covariance in $\mathrm{MRP\text{-}MLE}$.
We have shown in (10) that the asymptotic covariance of $\mathrm{MRP\text{-}MLE}$ goes from $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ when $n_{\mathrm{split}}=1$ to $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ when $n_{\mathrm{split}}\rightarrow\infty$. We have also established a qualified comparison in (9) that shows
$(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}\preceq(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}.$
However, $\bm{V}_{\mathrm{same}}^{\infty}$ and $\bm{V}_{\mathrm{diff}}^{\infty}$ are not explicit. Here we make a quantified illustration in a special case to better understand this shift in asymptotic covariance. We make a few simplifications to make it straightforward. First, we suppose that the user parameters $\zeta_{t}^{\star}$ are independently drawn from a distribution $\pi$, and we denote the expectation with respect to the random user parameter with $\mathbb{E}_{\mathrm{u}}$. Second, we set the item parameter to be $\bm{\theta}^{\star}=\bm{0}_{m}$. Third, we assume the sampling model where each user response to $mp$ items uniformly at random, for some even integer $mp$. We also need a constant $\beta$, which is a scalar defined by
$\beta\coloneqq\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{1}^{\star}}}{(e^{\zeta_{1}^{\star}}+1)^{2}}\right].$ (13)
Note that $\zeta_{1}^{\star}$ in this definition can be replaced by $\zeta_{t}^{\star}$ for any $t$.
In this special setting, we have the following result that quantifies the shift of asymptotic covariance for different $n_{\mathrm{split}}$. This proposition is special case for Theorem 3. The proof is deferred to Section D.3.
@@
Instate the assumptions of Theorem 1. For fixed $m$ and $p$, as $n\rightarrow\infty$,
$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\left(\frac{1}{n_{\mathrm{split}}}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\cdot\frac{mp}{2(mp-1)}\right)\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right);$ (14)
$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\cdot\frac{mp}{2(mp-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right).$
This proposition shows that in this special instance, the norm of the asymptotic covariance roughly scales as
$\frac{1}{n_{\mathrm{split}}}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\cdot\frac{mp}{2(mp-1)}.$
This equals $1$ when $n_{\mathrm{split}}=1$ and goes to mp/(2mp2)mp/(2mp-2) when $n_{\mathrm{split}}\rightarrow\infty$. The use of multiple random splitting in $\mathrm{MRP\text{-}MLE}$ or $\mathrm{WP\text{-}MLE}$ can shrink the asymptotic covariance of the estimators by a factor of mp/(2mp2)mp/(2mp-2).
4 Experiment
In this section, we demonstrate the empirical performance of $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ using both simulated and real data.
4.1 Simulations
We use simulated data to validate our theoretical results and compare our estimators with existing ones for the Rasch model. The data generating process follows the model specified in Section 2.1. Unless specified otherwise, in each trial, the user and item parameters are randomly drawn from
4.1.1 $\ell_{\infty}$ estimation error
We investigate the $\ell_{\infty}$ estimation error with the following goals:
  1. 1.
    We validate the theoretical result in $\ell_{\infty}$ estimation error of $\mathrm{RP\text{-}MLE}$ in Theorem 1.
  2. 2.
    We show how much advantage the $\mathrm{MRP\text{-}MLE}$ brings through multiple runs of data splitting.
  3. 3.
    We compare our methods with existing comparison-based algorithms, including the case where $\kappa_{1},\kappa_{2}$ are large.
Validating the theoretical result.
Theorem 1 tells us that the $\ell_{\infty}$ error scales as $1/\sqrt{np}$. Figure 1 shows that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ exhibits a near-linear relationship with respect to both $1/\sqrt{n}$ and $1/\sqrt{p}$, which is consistent with our theoretical predictions.
(a) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{n}$. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $10000$ to $40000$.
(b) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{p}$. The parameter is chosen to be $m=50,n=10000$ and $p$ varies from $1/9$ to $1$.
Figure 1: Estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ of $\mathrm{RP\text{-}MLE}$ with varying $n$ and $p$. Each point represents the average of 1000 trials.
Multiple runs in $\mathrm{MRP\text{-}MLE}$.
As we have discussed after Theorem 1, the random data splitting could incur a small loss of information. We have introduced a remedy $\mathrm{MRP\text{-}MLE}$ (Algorithm 2) to address this by averaging over multiple runs with independent data splitting. Moreover, in Proposition 5 we have a quantitative characterization of the improvement in $\ell_{2}$ error achieved through multiple data splittings.
Figure 2a shows that by averaging over more runs of data splittings, $\mathrm{MRP\text{-}MLE}$ achieves an improved $\ell_{\infty}$ estimation error that improves over PMLE and is close to $\mathrm{WP\text{-}MLE}$. In addition, we observe in Figure 2b that the improvement in squared $\ell_{2}$ error scales linearly with $1/n_{\mathrm{split}}$, consistent with the theoretical findings in Proposition 5.
(a) The $\ell_{\infty}$ estimation error of $\mathrm{MRP\text{-}MLE}$ v.s. $n_{\mathrm{split}}$. The dash-dotted and dashed lines are the performance of PMLE and $\mathrm{WP\text{-}MLE}$, respectively.
(b) The squared error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|^{2}$ of $\mathrm{MRP\text{-}MLE}$ v.s. $1/n_{\mathrm{split}}$ with varying $n_{\mathrm{split}}$. The dash-dotted and dashed lines are the performance of PMLE and $\mathrm{WP\text{-}MLE}$, respectively.
Figure 2: Estimation error of $\mathrm{MRP\text{-}MLE}$ with varying number of data splittings. For each trial, we record $\|\frac{1}{k}\sum_{i=1}^{k}\widehat{\bm{\theta}}_{(i)}-\bm{\theta}^{\star}\|$ for $k=1,\ldots,100$. The parameters are chosen to be $m=50,p=0.2,n=10000$. The latent scores are all 0 and the each user is assigned with mpmp item uniformly-at-random. Each point is averaged over 1000 trials.
Comparison with existing estimators.
We compare our algorithms with two other comparison-based algorithms: the pseudo MLE (PMLE) and the spectral method from NZ (23). In Figure 4, We can see that the performance of $\mathrm{MRP\text{-}MLE}$ is comparable to PMLE and slightly outperforms the spectral method. Our proposed algorithm not only offers stronger theoretical guarantees but also demonstrates competitive practical performance.
Performance with large $\kappa_{1},\kappa_{2}$.
While we assume $\kappa=\max\{\kappa_{1},\kappa_{2}\}=O(1)$ is most of this article, scenarios with large $\kappa$ can be practically relevant. To evaluate the performance in such cases, we compare the $\ell_{\infty}$ error of different methods under different condition numbers. For a fixed $\kappa$, we draw the user and item parameters as
and compute $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to have zero mean. Figure 4 illustrates the performance of different estimators as $\kappa$ varies. The MLE-based approaches including $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ achieve better $\ell_{\infty}$ error than the spectral method when $\kappa$ is large.
Figure 3: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $n$ using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$, and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $5000$ to $20000$. The result is averaged over 1000 trials.
Figure 4: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $\log(\kappa)$ using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$, and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1,n=20000$ and $\kappa$ varies from $1$ to $e^{10}$. The result is averaged over 1000 trials.
4.1.2 Top-$K$ recovery
We investigate the performance of different algorithms in top-$K$ recovery. Set $\theta_{i}^{\star}=(1-K/m)\Delta_{K}$ for $i\leq K$ and $\theta_{i}^{\star}=(-K/m)\Delta_{K}$ otherwise. For any estimator $\widehat{\bm{\theta}}$, we define top-$K$ recovery rate to be
$\frac{1}{K}\left|\left\{i\leq K:i\in A_{K}\right\}\right|,$
where $A_{K}$ is an arbitrary $K$-element set such that $\widehat{\theta}_{i}\geq\widehat{\theta}_{j}$ for any $i\in A_{K},j\notin A_{K}$. We compare the top-$K$ recovery rate of PMLE and the spectral method in NZ (23) with $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ in Figure 6. The recovery rate of PMLE, spectral method and $\mathrm{MRP\text{-}MLE}$ is similar, indicating again that our algorithm performs well in practice.
Figure 5: Top-$K$ recovery rate using spectral method, PMLE, $\mathrm{RP\text{-}MLE}$  and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=10000,m=50,p=0.1,K=5$ and $\Delta_{K}$ varies from 0.1 to 0.7. The result is averaged over 1000 trials.
4.1.3 Refined $\ell_{2}$ estimation error
In Theorem 4 we have shown that the $\ell_{2}$ error concentrate around $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ for $\widetilde{z}\in\left\{\widehat{z},z\right\}$. In each trial we compute the following quantity
[…]
4.1.4 Confidence intervals for $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$
Figure 7: Theoretical coverage rate $1-\alpha$ (dashed line) and empirical coverage rate $1-\widehat{\alpha}$ of the two-sided confidence intervals $\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$. The first 10 points on the left in each level are computed with $\mathrm{MRP\text{-}MLE}$ with $n_{\mathrm{split}}$ vaying from 10 to 100. The rightmost point on each level is computed with $\mathrm{WP\text{-}MLE}$. The parameters are set to be $n=10000,m=20,p=0.5$. The confidence interval is $1-\alpha=0.8,0.9,0.95,0.99$. Each point is averaged over $10000$ trials.
The asymptotic normality results in Theorem 3 and 4 allow us to construct confidence intervals for $\theta_{i}^{\star}$ with $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$. For large enough $n_{\mathrm{split}}$, we can approximate the variance of both estimators with $\frac{1}{n}(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}$, where $\bm{V}_{\mathrm{diff}}^{\infty}$ and $\bm{H}^{\infty}$ are estimated using the following plug-in estimates:
[…]
For a given confidence level $1-\alpha$, the confidence interval $\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$ is
$\mathcal{C}_{i}^{-}(\alpha/2)$ $\coloneqq\widehat{\theta}_{i}-z_{1-\alpha/2}\cdot\left[\frac{1}{n}(\widehat{\bm{H}}^{\infty})^{\dagger}\widehat{\bm{V}}_{\mathrm{diff}}^{\infty}(\widehat{\bm{H}}^{\infty})^{\dagger}\right]_{ii}^{1/2},$
$\mathcal{C}_{i}^{+}(\alpha/2)$ $\coloneqq\widehat{\theta}_{i}+z_{1-\alpha/2}\cdot\left[\frac{1}{n}(\widehat{\bm{H}}^{\infty})^{\dagger}\widehat{\bm{V}}_{\mathrm{diff}}^{\infty}(\widehat{\bm{H}}^{\infty})^{\dagger}\right]_{ii}^{1/2}.$
In Figure 7, we compare the empirical coverage rate
$1-\widehat{\alpha}\coloneqq\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\{\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\}.$
of these two-sided confidence intervals with the theoretical ones. We can see that when $n_{\mathrm{split}}$ is reasonably large, the empirical coverage rate matches well with the theoretical coverage rate for both $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$.
4.2 LSAT dataset
We study a real-world dataset (LSAT) on the Law School Admissions Test from DBL (70). LSAT has full observation of 1000 people answering 5 problems, with each person-item pair recording whether the answer was correct. The second row in Table 1 lists how many people answer each problem correctly. From the first look, Problem 3 appears to be the hardest question.
We proceed to quantify the hardness of these problems under the Rasch model and infer how confident we are in claiming it is the hardest. Using $\mathrm{MRP\text{-}MLE}$  we compute a latent score estimate and construct two-sided confidence intervals at significance level $\alpha=0.01$ for each coordinate, following the methodology introduced in Section 4.1.4. The result is summarized in Table 1, where higher latent score correspond to greater difficulty. The estimated parameters align inversely with the total number of correct answers. Notably, the lower bound of the confidence interval for $\theta_{3}^{\star}$ is larger than the upper bounds of the confidence intervals of $\theta_{1}^{\star},\theta_{2}^{\star},\theta_{4}^{\star}$, and $\theta_{5}^{\star}$. With Bonferroni correction, we can conclude that with 95% confidence Problem 3 is the most difficult problem in this dataset.
Problem 1 2 3 4 5
Total correct 924 709 553 763 870
$\theta$ estimate -1.2824 0.4511 1.2800 0.1926 -0.6413
CI lower bound -1.5579 0.2696 1.0958 0.0017 -0.8711
CI upper bound -1.0069 0.6327 1.4641 0.3834 -0.4116
Table 1: Latent score estimate calculated using $\mathrm{MRP\text{-}MLE}$ with 20 data splittings and confidence interval calculated with the construction introduced in Section 4.1.4. Higher latent score here means higher difficulty. The significance level is chosen to be $\alpha=0.01$ for each coordinate.
Now we assume Problem 3 is the top-$1$ item in latent score and investigate the top-$1$ recovery rate of different algorithms on LSAT under incomplete observation. In each trial we randomly select $\widetilde{n}$ people, and for each of them we randomly select their outcome on $\widetilde{m}$ problems. We then estimate $\bm{\theta}^{\star}$ using the subsampled data with different methods and compare the proportion of trials where the top-$1$ item is correctly identified. In Figure 8, we see results similar to the simulation. Our algorithm $\mathrm{MRP\text{-}MLE}$ has a similar recovery rate compared to PMLE and spectral method.
Figure 8: Top-1 recovery rate using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$  and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings. The parameters are chosen to be $\widetilde{m}=4$ and $\widetilde{n}$ varies from 100 to 200. The result is averaged over 10000 trials.
5 Discussion
In this paper, we propose two new likelihood-based estimators $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ for item parameter estimation in the Rasch model. Both enjoy optimal finite sample estimation guarantee and asymptotic normality that allows for tight uncertainty quantification. All this is achieved even when the user-item response data are extremely sparse (cf. NZ (23)). Below, we identify several questions that are interesting for further investigation:
  • Does PMLE or CMLE achieve optimal theoretical guarantee? In our experiments, pseudo MLE has shown a similar performance to $\mathrm{MRP\text{-}MLE}$. This naturally leads to the question of whether PMLE can enjoy the same theoretical guarantee. This is relevant to our work because our methods can be viewed as a modification of pseudo MLE by incorporating random disjoint pairing to decouple statistical dependency among paired $Y_{ij}$’s. It remains unclear whether such dependency is a fundamental bottleneck. On the other hand, conditional MLE is another popular method used in practice. In Section A.1 of the supplement we will mention that CMLE can also be viewed as a reduction to a less studied item-only model. This reduction is more complicated for analysis as it constructs item tuples rather than item pairs. It would be interesting to know whether we can transfer the techniques we have used here to develop a non-asymptotic analysis for CMLE.
  • Extending random pairing to other models in IRT. Some IRT models parameterize the latent score of users and items differently from the Rasch model. For instance, consider the two-parameter logistic model (2PL) with discrimination parameter on the users. It assumes that $X_{ti}$, the response of user $t$ to item $i$, follows the law
    $\mathbb{P}[X_{ti}=1]=\frac{1}{1+\exp(a_{t}^{\star}(\zeta_{t}^{\star}-\theta_{i}^{\star}))},$
    where $\bm{\theta}^{\star}$ is the latent scores of the items , $\bm{\zeta}^{\star}$ is the latent scores of the users, while $\bm{a}^{\star}$ is the discrimination parameters. Unlike the Rasch model, in the 2PL model,
    $\mathbb{P}[X_{ti}>X_{tj}\mid X_{ti}\neq X_{tj}]=\frac{\exp(a_{t}^{\star}\theta_{i}^{\star})}{\exp(a_{t}^{\star}\theta_{i}^{\star})+\exp(a_{t}^{\star}\theta_{j}^{\star})}$
    is not independent of the user discrimination parameter $\bm{a}^{\star}$. Therefore the reduction to the BTL model is no longer true in this case. However, one could employ a partially-Bayesian approach by putting a prior on $\bm{\alpha}^{\star}$ and maximize this marginally likelihood, which is a function of $\theta_{i}^{\star}-\theta_{j}^{\star}$ independent of $\bm{\zeta}$. It is interesting and non-trivial to extend the idea of random pairing to the 2PL model.
  • Extension to joint estimation of user and item parameters. It is sometimes of interest to estimate both the user and the item parameters. We expect our method $\mathrm{MRP\text{-}MLE}$ continues to work with slight modifications. In a high level, the idea is to estimate the mean-shifted parameters $\bm{\theta}^{\star}-(1/m)\bm{1}_{m}\bm{1}_{m}^{\top}\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}-(1/m)\bm{1}_{m}\bm{1}_{m}^{\top}\bm{\zeta}^{\star}$ using $\mathrm{MRP\text{-}MLE}$ twice. In the end, one estimates the difference in the means using MLE over the comparison outcomes. We leave the detailed investigation to future work.
Appendix A Analysis